# LLM-Mediated Guidance of MARL Systems

### v1

@@ 5 Experiments
To evaluate the effectiveness of RB and NL Controller interventions in our MARL framework, we conducted experiments within a custom AWS environment, part of the HIVEX suite. The experiments were designed to compare agents’ performance under three different intervention setups: No Controller, RB and NL Controller. For LLMs, we used Pharia-1-LLM-7B-control-aligned (AlephAlpha, 2024) or Llama-3.1-8B Instruct (Meta, 2023). Experiments assess how well intervention and non-intervention-supported agents can learn and perform. All experiment setups utilize Proximal Policy Optimization (PPO) as the MARL algorithm (Schulman et al., 2017) and are trained on $3\cdot 10^{5}$ time-steps. We use the default task (00) and terrain elevation level ($1$) of the AWS environment, but re-shaped rewards to focus on maximizing extinguishing tree rewards. We re-shaped the pick-up water reward from $1$ to $0.1$, the max preparing trees reward from $1$ to $0.1$ per tree, fire out reward from $10$ to 00, too close to village reward from $-50$ to 00, and the max extinguishing trees reward from $5$ to $1000$ per tree.
@@ 6 Results
Our results show that the RB and NL Controller interventions outperform the baseline without interventions, highlighting the potential of LLM-mediated guidance to accelerate training and enhance MARL performance in challenging environments. Generally, we can say that intervention is better than none, even with sparse supervision. In addition, both intervention controllers achieve a high-performance level and adapt to the demands of the new environment directive. Table 1 shows performance on extinguishing trees reward and episode mean reward for three controller setups: None, RB and NL for Pharia-1-7B-control-aligned and LLama-3.1-8B Instruct. In Figure 10, we show mean Extinguishing Trees Reward and in Figure 9 Episode Reward Mean over $10$ trials for each controller experiment, RB and NL versus the baseline without interventions. Please see Appendix A.7 for additional results.
We also investigated the scalability of our method by extending the default three-agent setup to
@@
Mediator Size Controller Episode R. Mean1 Ext. Trees R.2 Wall-Time3
None 238.34 (±14.34) 1.18 (±0.16) 2.65
Pharia-1-LLM 7B Rule-Based 437.65 (±43.28) 13.75 (±1.38) 2.96
Natural Language 372.05 (±24.45) 5.89 (±0.79) 3.988
Llama-3.1 8B Rule-Based 376.18 (±21.98) 15.76 (±1.76) 3.13
Natural Language 331.22 (±39.88) 6.73 (±0.81) 5.72
configurations with four, five, and six agents. Performance was compared between RB interventions and the no-intervention baseline using Episode Reward Mean and Extinguishing Trees Reward Mean for Pharia-1-7B-control-aligned and LLama-3.1-8B Instruct (Figure 11).
@@ 7 Discussion
A key observation is the comparative advantage of NL Controller interventions over non-intervention baselines. Pharia-1-LLM-7B-control-aligned outperformed in the Rule-Based Environment Mean Rewards, while Llama-3.1-8B Instruct excelled in the Extinguishing Trees Reward category. This suggests that Pharia-1-LLM-7B-control-aligned handles structured interventions better, while Llama-3.1-8B Instruct is more adept at free-form natural language interventions. The $300$-step intervention cooldown allowed agents to consolidate learning, operating independently for approximately $10$ steps. The adaptability of LLMs in real-time, context-sensitive guidance is evident, though each model excels in different dimensions. Both would benefit from memory of past tasks to refine strategies and enhance their adaptability in rapidly changing environments. The scalability experiments show that RB interventions consistently outperform the no-intervention baseline as agent numbers increase. Pharia-1 slightly outperforms LLama-3.1 in Episode Reward Mean, while both show a small decline in Extinguishing Trees Reward Mean with more agents, indicating coordination challenges in larger teams.
These findings suggest that LLM-based NL Controller interventions offer a promising approach for improving MARL systems, particularly where traditional RL methods face limitations. The distinct strengths of Pharia-1-LLM-7B-control-aligned and Llama-3.1-8B Instruct underscore the need for continued research to enhance LLM reasoning and planning capabilities. Further studies in more realistic environments are needed to validate these results across different domains.
@@ 8 Limitations and Potential Impacts
While our research demonstrates the significant potential of integrating LLMs into MA systems, several limitations and considerations must be acknowledged, particularly concerning bias, safety, the realism of the environment, and the transferability of our findings to other domains. Further discussion and information on resources and inference cost, and bias and safety concerns can be found in the Appendix A.1.
@@
Potential Impacts: Despite these limitations, the potential impacts of our research are substantial. By demonstrating the effectiveness of intervention and LLM-driven mediation in accelerating learning and improving coordination among agents, our approach offers a scalable solution for enhancing MARL systems in complex, dynamic environments. The findings suggest that human-like reasoning can lead to more efficient and effective learning processes, potentially reducing the computational resources required to train agents in complex environments. As these methods are refined and adapted to other domains, they could significantly advance the field of RL, contributing to more resilient and intelligent MA systems capable of tackling a wide range of real-world challenges.
@@ 9 Conclusion
In conclusion, our findings suggest that LLMs and MARL techniques have matured to a point where they can effectively adapt systems to complex, dynamic environments —an essential capability for tackling real-world challenges. The versatility of LLM-mediated interventions allows for easy adaptation to other domains, enabling efficient ’fine-tuning’ of MARL systems for specific tasks. While fully automating curriculum design remains challenging, minimal real-time human supervision can provide cost-effective, sparse guidance, helping agents develop more efficient policies and address increasingly complex tasks.
@@ Appendix A Appendix
A.1 Additional Limitations and Potential Impacts
Resources and Inference Cost: Another important consideration is the inference cost associated with the human LLM-mimicked interventions and LLM-Mediator. Out of the $3000$ total steps per agent per episode, the inference cost is only a fraction, as interventions are introduced every $300$ steps and typically influence agent behaviour for approximately $\sim 200$ steps. This periodic intervention minimizes the computational overhead, allowing agents to continue operating efficiently under the learned policy for the remaining $100$ steps. By balancing intervention frequency and task completion duration, we ensure that the computational load is manageable while still leveraging the benefits of real-time guidance from LLMs. Future work could further explore optimising this balance, reducing the task completion duration or intervention frequency while maintaining or improving agent performance. The training and testing of our experiment have been conducted on accessible, end-user hardware featuring an NVIDIA GeForce RTX $3090$ GPU, an AMD Ryzen $9$ $7950$X $16$-Core Processor, and $64$ GB of RAM. While these specifications align with high-end gaming laptops and desktop computers, the configuration could still be adapted to low-budget and non-GPU environments. This eliminates the need for specialized computational clusters, ensuring that researchers and practitioners with mid-range to high-end hardware can readily replicate our results using only consumer-grade equipment and an API for the LLM-Mediator.
A.2 Pseudocode
Standard PPO-CLIP pseudocode (OpenAI, 2021; Schulman et al., 2017):
Input: initial policy parameters $\theta_{0}$, initial value function parameters $\phi_{0}$
for $k=0,1,2,\ldots$ do
     Collect set of trajectories $\mathcal{D}_{k}$ = $\tau_{i}$} by running policy $\pi_{k}=\pi(\theta_{k})$ in the environment and
     overwriting with LLM-Mediator generated actions if an intervention has been issued.
     Compute rewards-to-go $\hat{R_{t}}$.
     Compute advantage estimates, $\hat{A_{t}}$ (using any method of advantage estimation) based on the
     current value function $V_{\phi_{k}}$
     Update the policy by maximizing the PPO-Clip objective:
[…]
     typically via stochastic gradient ascent with Adam.
     Fit value function by regression on mean-squared error:
     $\phi_{k+1}=arg\underset{\phi}{min}\frac{1}{|\mathcal{D}_{k}|T}\sum_{\tau\in\mathcal{D}_{k}}\sum_{t=0}^{T}\left((V_{\phi}(s_{t})-\hat{R_{t}}\right)$
     typically via some gradient descent algorithm.
end for
Algorithm 2
A.3 Hyperparameters
A.3.1 No Intervention
@@
  fire_close_to_city_reward: 0
@@
A.3.2 Rule-Based Llama-3.1-8B Instruct
@@
  fire_close_to_city_reward: 0
@@
A.3.3 Rule-Based Pharia-1-LLM-7B-control-aligned
@@
  fire_close_to_city_reward: 0
@@
A.3.4 Natural Language Llama-3.1-8B Instruct
@@
  fire_close_to_city_reward: 0
@@
A.3.5 Natural Language Pharia-1-LLM-7B-control-aligned
@@
  fire_close_to_city_reward: 0
@@ A.4 Environment Specification
  • Agent Count: $3$
  • Neighbour Count: 00
@@
  • Steer Left / Right (1): $[-1,1]$
@@
A.5 Un-shaped Reward Description and Calculation
@@
    Crossed Border - This is a negative reward of $-100$ given when the border of the environment is crossed. The border is a square around the island in the size of $1500$ by $1500$. The island is $1200$ by $1200$.
@@
    Pick-up Water - This is a positive reward of $1$ given when the agent steers the aeroplane towards the water. The island is $1200$ by $1200$ and there is a girdle of water around the island with a width of $300$.
@@
    Too Close to Village - This is a negative reward of $-50$ given when the fire is closer than $150$ to the centre of the village.
@@
    Time Step Burning - This is a negative reward of $-0.01$ given at each time-step, while the fire is burning.
@@ 2. Pick-up Water - To calculate the Pick-up Water reward, let us define the following:
    $r_{pw}=\begin{cases}1&\text{if }\left(p_{x}<eh\text{ or }p_{x}>-eh\text{ or }p_{y}<eh\text{ or }p_{y}>-eh\right)\\ &\text{and }\left(p_{x}>ih\text{ or }p_{x}<-ih\text{ or }p_{y}>ih\text{ or }p_{y}<-ih\right)\\ 0&\text{otherwise}\end{cases}$ (2)
@@ 3. Fire Out - To calculate the Fire Out reward, let us define the following:
    $r_{nb}=\begin{cases}10&\text{if }\forall t\in T,\,t\neq\text{"burning"}\\ 0&\text{otherwise}\\ \end{cases}$ (3)
@@ 5. Time Step Burning - To calculate the Time Step Burning reward, let us define the following:
  1. 1.
    We can now calculate the Time Step Burning reward:
    $r_{tsb}=\begin{cases}-0.01&\text{if }\forall t\in T,\,t=\text{"burning"}\\ 0&\text{otherwise}\\ \end{cases}$ (5)
@@ 6. Extinguishing Tree - To calculate the Extinguish Tree reward, let us define the following:
  1. 1.
    We can now calculate the Extinguish Tree reward:
[…]
@@ 7. Preparing Tree - To calculate the Preparing Tree reward, let us define the following:
  1. 1.
    We can now calculate the Preparing Tree reward:
[…]
A.6 Prompt Templates & Samples
A.6.1 Rule-Based Controller Prompt Template: LLM-Mediator
@@
A.6.2 Natural Language Controller Prompt Template: Strategy and LLM-Mediator
@@
A.6.3 Rule-Based and Natural Language Controller Vector Observation Data Samples
@@
A.6.4 Natural Language Strategy Samples: Pharia-1-LLM-7B-control-aligned
@@
A.6.5 Natural Language Strategy Samples: Llama-3.1-8B Instruct
@@ A.7 Additional Results
Figure 27: Fire too Close to City (Rule-Based) - No controller baseline VS Rule-Based Controller with Llama-3.1-8B Instruct: min, mean and max.
Figure 28: Fire too Close to City (Natural Language) - No controller baseline VS Natural Language Controller with Llama-3.1-8B Instruct: min, mean and max.

### v2

@@ 5 Experiments
To evaluate the effectiveness of RB and NL Controller interventions in our MARL framework, we conducted experiments within a custom AWS environment, part of the HIVEX suite. The experiments were designed to compare agents’ performance under three different intervention setups: No Controller, RB and NL Controller. For LLMs, we used Pharia-1-LLM-7B-control-aligned (AlephAlpha, 2024) or Llama-3.1-8B Instruct (Meta, 2023). Experiments assess how well intervention and non-intervention-supported agents can learn and perform. All experiment setups utilize Proximal Policy Optimization (PPO) as the MARL algorithm (Schulman et al., 2017) and are trained on $3\cdot 10^{5}$ time-steps. We use the default task (0) and terrain elevation level ($1$) of the AWS environment, but re-shaped rewards to focus on maximizing extinguishing tree rewards. We re-shaped the pick-up water reward from $1$ to $0.1$, the max preparing trees reward from $1$ to $0.1$ per tree, fire out reward from $10$ to 0, too close to village reward from 50-50 to 0, and the max extinguishing trees reward from $5$ to $1000$ per tree. Although the extinguishing reward is aggressively rescaled to emphasize the primary task objective, all controller and baseline experiments use identical reward functions, ensuring that observed performance differences are attributable to intervention strategies rather than reward design.
@@ 6 Results
Mediator Size Controller Episode R. Mean1 Ext. Trees R.2 Wall-Time3 None 238.34 (±14.34) 1.18 (±0.16) 2.65 Pharia-1-LLM 7B Rule-Based 437.65 (±43.28) 13.75 (±1.38) 2.96 Natural Language 372.05 (±24.45) 5.89 (±0.79) 3.99 Llama-3.1 8B Rule-Based 376.18 (±21.98) 15.76 (±1.76) 3.13 Natural Language 331.22 (±39.88) 6.73 (±0.81) 5.72
@@
Our results show that the RB and NL Controller interventions outperform the baseline without interventions, highlighting the potential of LLM-mediated guidance to accelerate training and enhance MARL performance in challenging environments. Generally, we can say that intervention is better than none, even with sparse supervision. In addition, both intervention controllers achieve a high-performance level and adapt to the demands of the new environment directive. Table 1 shows performance on extinguishing trees reward and episode mean reward for three controller setups: None, RB and NL for Pharia-1-7B-control-aligned and LLama-3.1-8B Instruct. In Figure 10, we show mean Extinguishing Trees Reward and in Figure 9 Episode Reward Mean over $10$ trials for each controller experiment, RB and NL versus the baseline without interventions. Please see Appendix A.5 for additional results. All reported results are averaged over 10 independent training runs with different random seeds; observed trends were consistent across runs.
@@
We also investigated the scalability of our method by extending the default three-agent setup to configurations with four, five, and six agents. Performance was compared between RB interventions and the no-intervention baseline using Episode Reward Mean and Extinguishing Trees Reward Mean for Pharia-1-7B-control-aligned and LLama-3.1-8B Instruct (Figure 11). As the number of agents increases, overall performance gradually decreases, indicating that coordination becomes more challenging in larger teams. However, this degradation is graceful, and LLM-mediated interventions consistently outperform the no-intervention baseline. Notably, the relative performance gap between intervention-based and non-intervention-based training widens with more agents, suggesting that centralized LLM guidance becomes increasingly beneficial as coordination complexity grows.
@@ 7 Discussion
A key observation is the comparative advantage of NL Controller interventions over non-intervention baselines. Pharia-1-LLM-7B-control-aligned outperformed in the Rule-Based Environment Mean Rewards, while Llama-3.1-8B Instruct excelled in the Extinguishing Trees Reward category. This suggests that Pharia-1-LLM-7B-control-aligned handles structured interventions better, while Llama-3.1-8B Instruct is more adept at free-form natural language interventions. The $200$-step intervention cooldown allowed agents to consolidate learning by operating independently between interventions. The adaptability of LLMs in real-time, context-sensitive guidance is evident. However, each model excels in different dimensions. Both would benefit from memory of past tasks to refine strategies and enhance their adaptability in rapidly changing environments. The scalability experiments show that RB interventions consistently outperform the no-intervention baseline as agent numbers increase. Pharia-1 slightly outperforms LLama-3.1 in Episode Reward Mean, while both show a small decline in Extinguishing Trees Reward Mean with more agents, indicating coordination challenges in larger teams.
These findings suggest that LLM-based NL Controller interventions offer a promising approach for improving MARL systems. This is particularly relevant in settings where traditional RL methods face fundamental limitations. The distinct strengths of Pharia-1-LLM-7B-control-aligned and Llama-3.1-8B Instruct underscore the need for continued research to enhance LLM reasoning and planning capabilities. Further studies in more realistic environments are needed to validate these results across different domains.
@@ 8 Limitations and Potential Impacts
While our research demonstrates the significant potential of integrating LLMs into MA systems, several limitations and considerations must be acknowledged, particularly concerning bias, safety, the realism of the environment, and the transferability of our findings to other domains.
@@
Resources and Inference Cost: Another important consideration is the inference cost associated with the human LLM-mimicked interventions and LLM-Mediator. Out of the $3000$ total steps per agent per episode, the inference cost is only a fraction, as interventions are introduced every $200$ steps and typically influence agent behaviour for approximately $\sim 120$ steps. This periodic intervention minimizes the computational overhead, allowing agents to continue operating efficiently under the learned policy for the remaining $\sim 80$ steps. By balancing intervention frequency and task completion duration, we ensure that the computational load is manageable while still leveraging the benefits of real-time guidance from LLMs. Future work could further explore optimising this balance, reducing the task completion duration or intervention frequency while maintaining or improving agent performance. The training and testing of our experiment have been conducted on accessible, end-user hardware featuring an NVIDIA GeForce RTX $3090$ GPU, an AMD Ryzen 99 79507950X $16$-Core Processor, and $64$ GB of RAM. While these specifications align with high-end gaming laptops and desktop computers, the configuration could still be adapted to low-budget and non-GPU environments. This eliminates the need for specialized computational clusters, ensuring that researchers and practitioners with mid-range to high-end hardware can readily replicate our results using only consumer-grade equipment and an API for the LLM-Mediator. Potential Impacts: Despite these limitations, the potential impacts of our research are substantial. By demonstrating the effectiveness of intervention and LLM-driven mediation in accelerating learning and improving coordination among agents, our approach offers a scalable solution for enhancing MARL systems in complex, dynamic environments. The findings suggest that human-like reasoning can lead to more efficient and effective learning processes, potentially reducing the computational resources required to train agents in complex environments. As these methods are refined and adapted to other domains, they could significantly advance the field of RL, contributing to more resilient and intelligent MA systems capable of tackling a wide range of real-world challenges.
@@ 9 Conclusion
In conclusion, our findings suggest that LLMs and MARL techniques have matured to a point where they can effectively adapt systems to complex, dynamic environments – an essential capability for tackling real-world challenges. The versatility of LLM-mediated interventions allows for easy adaptation to other domains, enabling efficient training of MARL systems for specific tasks. While fully automating curriculum design remains challenging, minimal real-time human supervision can provide cost-effective, sparse guidance, helping agents develop more efficient policies and address increasingly complex tasks.
@@ Appendix A Appendix
A.1 Hyperparameters
A.1.1 No Intervention
@@
  fire_close_to_village_reward: 0
@@
A.1.2 Rule-Based Llama-3.1-8B Instruct
@@
  fire_close_to_village_reward: 0
@@
A.1.3 Rule-Based Pharia-1-LLM-7B-control-aligned
@@
  fire_close_to_village_reward: 0
@@
A.1.4 Natural Language Llama-3.1-8B Instruct
@@
  fire_close_to_village_reward: 0
@@
A.1.5 Natural Language Pharia-1-LLM-7B-control-aligned
@@
  fire_close_to_village_reward: 0
@@ A.2 Environment Specification
  • Agent Count: 33
  • Neighbour Count: 0
@@
  • Steer Left / Right (1): [1,1][-1,1]
@@
A.3 Un-shaped Reward Description and Calculation
@@
    Crossed Border - This is a negative reward of 100-100 given when the border of the environment is crossed. The border is a square around the island in the size of $1500$ by $1500$. The island is $1200$ by $1200$.
@@
    Pick-up Water - This is a positive reward of $1$ given when the agent steers the aeroplane towards the water. The island is $1200$ by $1200$ and there is a surrounding band of water around the island with a width of $300$.
@@
    Too Close to Village - This is a negative reward of 50-50 given when the fire is closer than $150$ to the centre of the village.
@@
    Time Step Burning - This is a negative reward of 0.01-0.01 given at each time-step, while the fire is burning.
@@ 2. Pick-up Water - To calculate the Pick-up Water reward, let us define the following:
[…]
@@ 3. Fire Out - To calculate the Fire Out reward, let us define the following:
[…]
@@ 5. Time Step Burning - To calculate the Time Step Burning reward, let us define the following:
$r_{tsb}=\begin{cases}-0.01&\text{if }\exists t\in T\text{ such that }t=\text{``burning''}\\ 0&\text{otherwise}\end{cases}$ (5)
@@ 6. Extinguishing Tree - To calculate the Extinguish Tree reward, let us define the following:
$r_{e}=5\sum_{t\in T}\mathbb{I}\!\left(t_{\text{previous}}=\text{``burning''}\;\land\;t_{\text{current}}=\text{``extinguished''}\right)$ (6)
@@ 7. Preparing Tree - To calculate the Preparing Tree reward, let us define the following:
$r_{p}=\sum_{t\in T}\mathbb{I}\!\left(t_{\text{previous}}=\text{``not burning''}\;\land\;t_{\text{current}}=\text{``wet''}\right)$ (7)
A.4 Prompt Templates & Samples
A.4.1 Rule-Based Controller Prompt Template: LLM-Mediator
@@
A.4.2 Natural Language Controller Prompt Template: Strategy and LLM-Mediator
@@
A.4.3 Rule-Based and Natural Language Controller Vector Observation Data Samples
@@
A.4.4 Natural Language Strategy Samples: Pharia-1-LLM-7B-control-aligned
@@
A.4.5 Natural Language Strategy Samples: Llama-3.1-8B Instruct
@@ A.5 Additional Results
Figure 27: Fire too Close to Village (Rule-Based) - No controller baseline VS Rule-Based Controller with Llama-3.1-8B Instruct: min, mean and max.
Figure 28: Fire too Close to Village (Natural Language) - No controller baseline VS Natural Language Controller with Llama-3.1-8B Instruct: min, mean and max.